Subject: Interim 3.5-m telescope report
From: Bruce Gillespie
Submitted: Fri, 24 Aug 2007 13:08:18 -0600
Message number: 1050
(previous: 1049,
next: 1051
up: Index)
All:
Our tradition is to skip one of the Users Committee telecons during
the summer, which for this year is the August meeting. In lieu of
our usual meeting minutes, Suzanne has asked me to send you an
interim telescope, instruments, and engineering report on the 3.5-m
telescope, prepared by Mark Klaene.
bg
**************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 7/6/07
through 8/17/07
Mark Klaene
0) Overview
Monsoonal weather continued with moderate rain and evening clouds,
limiting observations for much of July and early August. The
engineering time did yield some on-sky time with useful data.
1) Telescope
Tertiary rotation problems showed up causing a lot of work to locate
a culprit. It was finally determined that a loose wire along with
the Galil controller was causing the problems. Tracking tuning was
improved.
The summer shutdown will start August 20. The task list is included
below. It is an ambitious schedule with very limited days to check
items on the sky so weather could easily cause a slippage in
returning to science.
2) Instruments
Agile commissioning was completed with the focal reducer and fiber
interface. We are reviewing the data and the plans to incorporate it
on a mirror cell side port.
The replacement DIS Red high-resolution grating has been installed,
aligned and tested and placed in service. We will look at the issue
of the spectra appearing near the bottom of the detector during the
shutdown.
The Lakeshore temperature controller was repaired and placed back in
service.
Incorporation of SPIcam into TUI is completed and will be the default
for users when we return from shutdown.
NIC-FPS was taken off line and sent to Boulder on the 26th of July.
The etalon is running in the test dewar, cold; the engineering chip
is now in NIC-FPS for etalon testing. The science-grade chip has
been returned from Rockwell with the 16-channel upgrade. Everything
is on track for a return of NIC-FPS to APO in mid-September.
4) Miscellaneous
Jim Davenport completed his summer fill-in as an Observing Specialist
and will now be moving on to graduate school. Jim's help was
invaluable this summer, he did a fantastic job operating the
telescope and his insights and software work proved very useful.
Summer Shutdown Task List
M3
Re-aluminize
Service actuators
M1
Re-aluminize
Check, photograph cracks
PMSS
Pressure transducer calibration
Load cell calibration
Bellofram inspection leak check
Air line leak check with belloframs by sector
Servo valves inspection/cleaning
Clean mirror cell
Vent tube repair
Foam inspection
Mirror cover actuator insp/service
Calibrate LVDTs
Investigate fore/aft mirror constraint
Mirror covers
Re-paint white
Repair rivets
Pump SPIcam
Pump DIS
Purge/recharge blue camera cryotiger system
Install new bushings in grating holders
Remount DIS red high grating in new holder
Replace UPS batteries
Echelle
Inst - I level
Elect I -level
#3 for ops bldg
Echelle electronics fan replacement
New axis controller for rotator
New NA2 rotator assembly
Repair strain gauge metrology noise issue
Replace temp sensors on primary mirror
Repair 2ndary accelerometer recording system
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1050 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO